### v1

# Energy dependence of cross sections in proton–proton and antiproton–proton collisions

Abstract
Abstract—Energy dependence of global scattering parameters, mostly of total cross section, is studied for proton–proton and antiproton–proton collisions. Results are presented for physical analysis updated with taken into account the recent data from accelerator experiments as well as from cosmic ray measurements. The analytic parameterizations suggested within Axiomatic Quantum Field Theory (AQFT) provide the quantitative description of energy dependence of global scattering parameters for rather wide energy range. Detailed scan on low boundary of the fitting range for energy dependence of global scattering parameters allows the observation of the onsets for regions in which Pomeranchuk theorem and / or Froissart–Martin one is valid. It is obtained that global scattering parameters show the behavior corresponded to any formulations of Pomeranchuk theorem and closed to (modified) Froissart–Martin limit in functional sense in multi-TeV energy region. Bosonic condensation is considered as one of the possible dynamical mechanisms which would be provide the total cross section approaches to (modified) Froissart–Martin limit at quantitative level but not functionally only.
pacs:
13.85.Dz, 13.85.-t
I Introduction
The large-scale projects of hadronic colliders of next generation assume the measurements for proton–proton ($pp$) interactions at collision energies up to $\sqrt{s}\sim 100$ TeV in order to magnitude arXiv-1507.03224-2015 ; EPJC-79-474-2019 . Phenomenological study of cross sections as one the most important global quantities for (anti)proton–proton ($\bar{p}p$, pppp) scattering at very high $\sqrt{s}$ can be helpful for search for the onset of the asymptotic regime in energy dependence of cross sections and related parameters as well as for the establish of more precision and reliable criteria for asymptotic domain in both mathematical and physical senses. The rigorous theorems and relations between global scattering parameters are crucially important for high energy physics, particularly for the asymptotic energy domain. In this paper, two of the most outstanding among them, namely the Pomeranchuk theorem JEPT-7-499-1958 and the Froissart or, perhaps, more correctly Froissart–Martin one PR-123-1053-1961 ; PR-129-1432-1963 are considered, both concerning $pp$ and $\bar{p}p$ scattering processes. Importance of the asymptotic theorems for high energy physics is in the ability for direct and model-independent verification of a most general and basic principles which are base for Quantum Field Theory (QFT) at all. However, the theory, while predicting the behavior of some global scattering parameters for particle–particle and antiparticle–particle scattering processes, provides no indications and / or estimates, even by order of magnitude, of the energies at which the asymptotic region begins. As it is known a limit in the mathematical sense is only indicated. Various observables and its combinations are used for the study of asymptotic regime Okorokov-IJMPA-33-1850077-2018 ; Okorokov-PAN-81-508-2018 ; Okorokov-PAN-82-134-2019 ; Okorokov-CPC-46-083105-2022 . Estimations for low boundary on $\sqrt{s}$ for onset of asymptotic regime depend on model and / or choice of a global observable used for consideration. Such estimations can differ (very) significantly from each other. Furthermore the definition of term ”asymptote” itself in physical sense for global scattering parameters or ”truly asymptotic region” is non-trivial Okorokov-IJMPA-33-1850077-2018 . As consequence, the study of global scattering parameters at high and very high collision energies, in particular, for direct rigorous verification of asymptotic theorem is rather complex problem in various sub-fields (theory, phenomenology, data analysis, experiment) of modern physics of strong interaction. Therefore, any estimates, even qualitative ones in order of magnitude, of the lower energy bound for the onset of the asymptotic region seem important. Such information can be obtained using phenomenological models and physical analysis of experimental data. Perhaps, such studies may be actual both for a better understanding of the dynamics of strong interactions at large distances (small-angle scattering) in wide energy domain and for the development of physics research programs for future accelerator facilities. In the paper the Axiomatic Quantum Field Theory (AQFT) approach is used for phenomenological study of energy dependence of global scattering parameters, mostly, of total cross sections.
II Asymptotic relations and AQFT model
The Pomeranchuk theorem allows the two formulations Okorokov-CPC-46-083105-2022 . The original version assumes that if the forward elastic scattering amplitude $\mathcal{F}^{\pm}(s)$, where the plus sign is for $pp$ and the minus one for $\bar{p}p$ interaction [^1] As indicated above, the theorem is valid for general case particle–particle and antiparticle–particle processes but here the formulation is adopted for specific reactions under study. for brevity, have no oscillations at $s\to\infty$ and they are mostly imaginary $\lim_{\varepsilon\to\infty}[\mathrm{Re}\mathcal{F}^{\pm}]/[\mathrm{Im}\mathcal{F}^{\pm}\ln\varepsilon]=0$ with $\varepsilon\equiv s/s_{0}$, $s_{0}=1$ GeV2, then the difference between $pp$ and $\bar{p}p$ total cross sections tend to zero in asymptotic region
$\Delta_{\footnotesize\mbox{t}}(s)\equiv|\sigma_{\scriptsize{\mbox{t}}}^{+}-\sigma_{\scriptsize{\mbox{t}}}^{-}|\rightarrow 0,~\mathrm{at}~s\to\infty.$ (1)
This relation was derived under the assumption of asymptotically constant cross sections $\sigma_{\scriptsize{\mbox{t}}}^{\pm}\to C^{\pm}_{\scriptsize{\mbox{P}}}$ at $s\to\infty$, where $C^{\pm}_{\scriptsize{\mbox{P}}}\neq 0$ are some constants and $|C^{+}_{\scriptsize{\mbox{P}}}-C^{-}_{\scriptsize{\mbox{P}}}|\to 0$ at $s\to\infty$ as the condition for correctness of (1).
The modified, or generalized for increasing cross sections, formulation of the theorem means that
$R_{\footnotesize\mbox{t}}^{-/+}(s)\equiv\sigma_{\footnotesize\mbox{t}}^{-}(s)/\sigma_{\footnotesize\mbox{t}}^{+}(s)\rightarrow 1,~\mathrm{at}~s\rightarrow\infty.$ (2)
This asymptote was proved if $\sigma_{\scriptsize{\mbox{t}}}^{\pm}$ grow with energy and this behaviour of $\sigma_{\scriptsize{\mbox{t}}}^{\pm}(s)$ is just observed in experimentally attainable high-energy domain.
As it is known the Froissart–Martin theorem is formulated as inequality (upper bound) and it states that any total cross section cannot grow faster than $\ln^{2}\varepsilon$. There are two various views of inequality within the Froissart–Martin theorem in different publications [^2] Future discussion will demonstrate that this discrepancy in the formulation of Froissart–Martin theorem will not affect on the physical conclusions based on the available experimental databases.. It is possible the conditional inequality Bogolubov-book-1990-1 and absolute one PRD-80-065013-2009 ; the corresponding mathematical views of Froissart–Martin theorem are
$\sigma_{\scriptsize{\mbox{t}}}\leq\sigma_{\scriptsize{\mbox{FM}}},~\mathrm{at}~s\to\infty;$ (3a)
$\sigma_{\scriptsize{\mbox{t}}}<\sigma_{\scriptsize{\mbox{FM}}},~\mathrm{at}~s\to\infty;$ (3b)
According to the study PRD-80-065013-2009 the new bound from Froissart–Martin theorem or, equivalently, modified Froissart–Martin limit is $\sigma_{\scriptsize{\mbox{FM}}}=C^{(m)}_{\scriptsize{\mbox{FM}}}\ln^{2}\varepsilon$, $C^{(m)}_{\scriptsize{\mbox{FM}}}\equiv\pi/2m^{2}_{\pi}\approx 31.4$ mb and $m_{\pi}$ is the pion mass PRD-110-030001-2024 . The modified limit is twice smaller than that from original studies NCA-42-930-1966 based on the principles of AQFT Okorokov-CPC-46-083105-2022 due to relation between constant terms $C_{\scriptsize{\mbox{FM}}}=2\,C^{(m)}_{\scriptsize{\mbox{FM}}}$ for original and modified limits. The (modified) limit does not depend on the type of collisions.
The theorems of Pomeranchuk and Froissart–Martin are independent from each other mathematically and physically. In general, the energy domains can be different for the validity of each of these asymptotic theorems. Therefore, the more correct and careful statement is the reach of the energy domain for validity of certain theorem or asymptotic region / regime for fixed observable with point of view of certain theorem if any. But similar observation will remain the problem of reach of ”truly asymptotic region” still open and without rigorous quantitative answer.
The analytic parameterizations were proposed within AQFT Okorokov-IJMPA-25-5333-2010 in order to describe the energy dependence of the most known global scattering parameters, namely, $\sigma_{\scriptsize{\mbox{t}}}^{\pm}$ and $\rho$-parameter defined as the ratio $\rho^{\pm}=\mathrm{Re}\mathcal{F}^{\pm}/\mathrm{Im}\mathcal{F}^{\pm}$ in $pp$ and $\bar{p}p$ scattering. The following form of the parameterizations
$\sigma_{\scriptsize{\mbox{t}}}^{\pm}(s)$ $=$ $\frac{\delta_{\pm}}{\varepsilon^{6+\gamma_{\pm}}}+\frac{\beta_{\pm}}{\varepsilon^{\alpha_{\pm}}}\,\sigma_{\scriptsize{\mbox{FM}}},$ (4)
$\rho^{\pm}(s)$ $=$ $\frac{1}{\sigma_{\scriptsize{\mbox{t}}}^{\pm}}\biggl\{\frac{k}{\varepsilon}+\frac{\pi}{2}\biggl[-\frac{\delta_{\pm}(6+\gamma_{\pm})}{\varepsilon^{6+\gamma_{\pm}}}+\frac{\beta_{\pm}}{\varepsilon^{\alpha_{\pm}}}\,\sigma_{\scriptsize{\mbox{FM}}}\biggl(\frac{2}{\ln\varepsilon}-\alpha_{\pm}\biggr)\biggr]\pm\frac{\pi}{4}[\sigma_{\footnotesize\mbox{t}}^{+}-\sigma_{\footnotesize\mbox{t}}^{-}]\biggr\}$ (5)
is more useful for the present study. Here the factor $\sigma_{\scriptsize{\mbox{FM}}}$ corresponding to the modified Froissart–Martin limit PRD-80-065013-2009 is singled out in clear view.
The relations between free parameters are shown in Table 1 at which various formulations of Pomeranchuk theorem are valid within AQFT approach. In most part of cases formulations (1), (2) do not allow numerical estimations of parameters $\alpha$, $\beta$, $\gamma$, $\delta$ at which Pomeranchuk theorem is valid.
Table 1: Relations for AQFT parameters for validity of Pomeranchuk theorem.
$\alpha$ $\beta$ $\gamma$ $\delta$
Original formulation
$\alpha_{+}=\alpha_{-}\equiv\alpha_{\scriptsize{\mbox{P}}}$ $\beta_{+}=\beta_{-}\equiv\beta_{\scriptsize{\mbox{P}}}$ $\gamma_{+}=\gamma_{-}\equiv\gamma_{\scriptsize{\mbox{P}}}>-6$ $\delta_{\pm}$ are finite
$\gamma_{+}=\gamma_{-}\equiv\gamma_{\scriptsize{\mbox{P}}}=-6$ $\delta_{+}=\delta_{-}\equiv\delta_{\scriptsize{\mbox{P}}}$
Generalized formulation
$\alpha_{+}=\alpha_{-}\equiv\alpha_{\scriptsize{\mbox{P}}}$ $\beta_{+}=\beta_{-}\equiv\beta_{\scriptsize{\mbox{P}}}$ $\gamma_{+}=\gamma_{-}\equiv\gamma_{\scriptsize{\mbox{P}}}>-6$, $\delta_{\pm}$ are finite
$\gamma_{\scriptsize{\mbox{P}}}+6\geq\alpha_{\scriptsize{\mbox{P}}}$
$\beta_{\pm}$ are finite $\gamma_{+}=\gamma_{-}\equiv\gamma_{\scriptsize{\mbox{P}}}=-6$ $\delta_{+}=\delta_{-}\equiv\delta_{\scriptsize{\mbox{P}}}$
The observables $\sigma_{\scriptsize{\mbox{t}}}^{\pm}$ are characterized by the one energy dependence (4) in functional sense. Thus the (modified) Froissart–Martin limit is reached at the equal values of AQFT parameters for $pp$ and $\bar{p}p$ collisions. These asymptotic values with point of view of the Froissart–Martin theorem are shown in Table 2.
Table 2: Values of AQFT parameters for modified Froissart–Martin limit.
$\alpha_{\scriptsize{\mbox{FM}}}$ $\beta_{\scriptsize{\mbox{FM}}}$ $\gamma_{\scriptsize{\mbox{FM}}}$ $\delta_{\scriptsize{\mbox{FM}}}$
0.0 1.0 6.0 0.0
It is seen from Tables 1, 2 the reach of the (modified) Froissart–Martin limit by $\sigma_{\scriptsize{\mbox{t}}}^{\pm}$ leads to the validity of Pomeranchuk theorem in any formulations within AQFT. But, in general, the opposite statement is not correct.
The asymptotic theorems considered here establish some relations for total cross sections but they do not provide any additional information and / or asymptotic relations for $\rho$-parameter. From the request of mostly imaginary $\mathcal{F}^{\pm}(s)$ at $s\to\infty$ it only consequents $\rho^{\pm}=\mathrm{Re}\mathcal{F}^{\pm}/\mathrm{Im}\mathcal{F}^{\pm}\ll 1$ in the energy domain of validity of Pomeranchuk theorem. Within AQFT approach the $\rho$-parameter as well as total cross section is equal for $pp$ and $\bar{p}p$ collisions in the energy domain of (modified) Froissart–Martin limit and $\rho$ reaches the asymptotic value $\rho_{\scriptsize{\mbox{FM}}}=\pi\ln^{-1}\varepsilon$ because $\kappa/\varepsilon\to 0$ at $\varepsilon\to\infty$ at any finite $\kappa$. That asymptotic level for $\rho$ agrees with results for $\left.\rho(s)\right|_{s\to\infty}$ obtained with help of the crossing property, 1st order derivative dispersion relations and optical theorem Okorokov-CPC-46-083105-2022 .
III Experimental data and fit procedure
The main global physical quantities characterizing the elastic scattering process of two particles are the well-known total cross section ($\sigma_{\footnotesize\mbox{t}}$) and the ratio of the real to the imaginary part of the forward scattering amplitude ($\rho$ parameter). Therefore the set of global parameters $\mathcal{G}_{1}=\{\mathcal{G}_{1}^{i}\}_{i=1}^{4}=\{\sigma_{\scriptsize{\mbox{t}}}^{+},\sigma_{\scriptsize{\mbox{t}}}^{-},\rho^{+},\rho^{-}\}$ is under study for elastic $pp$ and $\bar{p}p$ scattering below. The set $\mathcal{G}_{1}$ contains only observables that are independent of each other as well as directly measured in experiments.
The choice of the ”standard” database for the present work was described in details elsewhere Okorokov-IJMPA-25-5333-2010 ; Okorokov-IJMPA-32-1750175-2017 . The experimental database for $\mathcal{G}_{\scriptsize{1}}$ contained the ensembles for $\sigma_{\scriptsize\mbox{t}}^{\pm}$, $\rho^{\pm}$ from PRD-110-030001-2024 is standard one. That database is denoted as DB24 while the database taken into account the above samples and results for $\sigma_{\scriptsize\mbox{t}}^{+}$ at $\sqrt{s}=0.20$, 13 TeV from STAR PLB-808-135663-2020 and ATLAS EPJC-83-441-2023 respectively, for $\rho^{+}$ at $\sqrt{s}=13$ TeV from ATLAS EPJC-83-441-2023 and TOTEM EPJC-79-785-2019 is referred as DB24+. Influence of the TOTEM results for $\rho^{+}$ on the version of DB24+ was analyzed in details elsewhere Okorokov-CPC-46-083105-2022 . Table 3 summarizes the main features of the databases of experimental results discussed in the present work for the set of the scattering parameters $\mathcal{G}_{\scriptsize{1}}$.
The set of free parameters used within AQFT approach $\mathcal{P}=\{\mathcal{P}^{i}\}_{i=1}^{9}=\{\alpha_{+,-},\beta_{+,-},\gamma_{+,-},\delta_{+,-},\kappa\}$ can be obtained from the fitting procedures of experimental data. One can stress that each of $\sigma_{\scriptsize{\mbox{t}}}^{\pm}$ is driven by the corresponding subset $\mathcal{P}_{\pm}=\{\mathcal{P}^{i}_{\pm}\}_{i=1}^{4}=\{\alpha_{\pm},\beta_{\pm},\delta_{\pm},\gamma_{\pm}\}$ of 4 free parameters whereas $\rho^{\pm}$ depends on full set $\mathcal{P}$ of 9 free parameters due to interference term $\sigma_{\footnotesize\mbox{t}}^{+}-\sigma_{\footnotesize\mbox{t}}^{-}]$ in (5). Thus, experimental data for $\rho$-parameter for only one ($pp/\bar{p}p$) scattering allow, in general, the derivation of values for full set $\mathcal{P}$ of free parameters, i.e. for both $pp$ and $\bar{p}p$ reactions, within AQFT approach.
The energy dependence of the terms of $\mathcal{G}_{\scriptsize{1}}$ formed from DB24 is approximated by (4), (5) in the range $\sqrt{s}\geq\sqrt{s_{\scriptsize{\mbox{min}}}}$, where $\sqrt{s_{\scriptsize{\mbox{min}}}}$ is the low boundary of fit range which is defined empirically. In the present work the wide collection $\bigl\{\sqrt{s_{\scriptsize{\mbox{min}},i}}\bigr\}_{i=1}^{16}=\{3,5,10,15,20,25,30,40,50,60,86.45,100,500,10^{3},5\times 10^{3},10^{4}\}$ GeV is used in order to detailed scan $\sigma_{\scriptsize{\mbox{t}}}^{\pm}(s)$, $\rho^{\pm}(s)$ and search for possible approach to the asymptotic behavior [^3] The set $\bigl\{\sqrt{s_{\scriptsize{\mbox{min}},i}}\bigr\}_{i=1}^{16}$ considered here is the joined one for studies Okorokov-CPC-46-083105-2022 , Okorokov-IJMPA-25-5333-2010 ; Okorokov-IJMPA-32-1750175-2017 and the point $\sqrt{s_{\scriptsize{\mbox{min}},11}}=86.45$ GeV corresponded the appearance of the first data point from cosmic-ray measurements for $\sigma_{\scriptsize{\mbox{t}}}^{+}$ Okorokov-PAN-81-508-2018 .. For the respective experimental points, the total uncertainties are considered which are calculated as $\Delta^{2}_{\scriptsize{\mbox{tot}}}=\Delta^{2}_{\scriptsize{\mbox{stat}}}+\Delta^{2}_{\scriptsize{\mbox{sys}}}$, where $\Delta^{2}_{\scriptsize{\mbox{stat (sys)}}}$ stands for statistical (systematic) errors.
Table 3: Databases for the set of global scattering parameters $\mathcal{G}_{\scriptsize{1}}$.
Database Parameter from the set $\mathcal{G}_{1}$
$\sigma_{\scriptsize\mbox{t}}^{+}$ $\sigma_{\scriptsize\mbox{t}}^{-}$ $\rho^{+}$ $\rho^{-}$
DB24    PRD-110-030001-2024    PRD-110-030001-2024    PRD-110-030001-2024    PRD-110-030001-2024
DB241+    PRD-110-030001-2024 ; PLB-808-135663-2020 ; EPJC-83-441-2023    –//–    PRD-110-030001-2024 ; EPJC-83-441-2023 and $\left.\langle\rho^{pp}\rangle\right|_{\sqrt{s}=13\,\scriptsize{\mbox{TeV}}}$ Okorokov-CPC-46-083105-2022    –//–
DB202+    –//–    –//–    PRD-110-030001-2024 ; EPJC-79-785-2019 ; EPJC-83-441-2023    –//–
It should be stressed that there are only few experimental points for $\bar{p}p$ reaction at $\sqrt{s}>1$ TeV. Moreover the data for that interaction are absent at $\sqrt{s}>5$ TeV at all. Therefore the simultaneous fit for global scattering parameters is made for $pp$ only at $\sqrt{s_{\scriptsize{\mbox{min}},15}}=5$ TeV and numerical values of terms of the full set $\mathcal{P}$ are derived due to (5) for $\rho^{+}(s)$. In DB24 there are experimental results for $\sigma_{\footnotesize\mbox{t}}^{+}(s)$ only at $\sqrt{s_{\scriptsize{\mbox{min}},16}}=10$ TeV and, consequently, simultaneous fit procedure reduces to the individual fit for $pp$ total cross section by (4). Similar study was made for AQFT in our previous work Okorokov-PAN-81-508-2018 . Here the individual fit for $\sigma_{\footnotesize\mbox{t}}^{+}(s)$ is not considered. The simultaneous fit of $pp$ part of the set $\mathcal{G}_{1}$$\mathcal{G}_{+}=\{\mathcal{G}_{+}^{j}\}_{j=1}^{2}=\{\sigma_{\scriptsize{\mbox{t}}}^{+},\rho^{+}\}$ – can be made for versions of DB24+ (Table 3) at highest $\sqrt{s_{\scriptsize{\mbox{min}},16}}=10$ TeV. This work is in the progress. The features of DB24 and fit procedure at high $\sqrt{s_{\scriptsize{\mbox{min}}}}\geq 1$ TeV, in particular available $\bar{p}p$ data, should be taken into account for future discussion.
IV Fit results and discussion
In the present work simultaneous fits are made by equations (4), (5) for the energy dependence of global scattering parameters from the set $\mathcal{G}_{1}$ for DB24. The numerical results of the simultaneous fits are shown in Table 4 for low boundaries $s_{\scriptsize{\mbox{min}},i}$, i=115i=1-15 of approximated energy range [^4] The first value in the cell of Table 4 corresponds to the model parameter with plus sign, i.e. for $pp$, and second value – to the model parameter with minus sign, i.e. for $\bar{p}p$; $\chi^{2}_{1}\equiv\chi^{2}$/n.d.f.. Experimental data from DB24 together with fit curves are shown in Fig. 1. Also additional points from DB242+ points are shown for completeness, but as indicated above the points from DB24 are only fitted in Fig. 1. Here solid line correspond to the fits at $\sqrt{s_{\scriptsize{\mbox{min}},2}}=5$ GeV, dashed curves are fits at $\sqrt{s_{\scriptsize{\mbox{min}},7}}=30$ GeV and dotted lines are for $\sqrt{s_{\scriptsize{\mbox{min}},14}}=1$ TeV.
Table 4: Results for simultaneous fitting by equations (4), (5) of $\mathcal{G}_{1}(s)$ for DB24.
$i$ Parameter
$\alpha_{\pm}\times 10^{3}$ $\beta_{\pm}\times 10^{3}$ $\gamma_{\pm}\times 10^{3}$ $\delta_{\pm}$, mb $\kappa$, mb $\chi^{2}_{1}$
1 $20.0\pm 1.3$ $12.84\pm 0.29$ $5886.1\pm 2.0$ $51.51\pm 0.22$ $42\pm 5$ 1.88
$76.4\pm 2.1$ $32.0\pm 0.8$ $5682\pm 4$ $124.9\pm 1.4$
2 $18.1\pm 2.3$ $12.4\pm 0.6$ $5890\pm 5$ $51.0\pm 0.7$ $46\pm 14$ 1.13
$52\pm 4$ $22.4\pm 1.2$ $5763\pm 8$ $96.0\pm 2.5$
3 $6\pm 4 11$$1.9\pm 1.1$ $5895\pm 11$ $50.3\pm 1.9$ $110\pm 30$ 1.03
$36\pm 7$ $17.3\pm 2.0$ $5815\pm 18$ $79\pm 5$
4 $4\pm 6 11$$1.4\pm 1.3$ $5900\pm 15$ $49.6\pm 2.7$ $42\pm 29$ 1.00
$20\pm 9$ $12.8\pm 2.1$ $5864\pm 20$ $65\pm 5$
5 $22\pm 6$ $13.6\pm 1.7$ $873\pm 21 55$$5\pm 5$ $35\pm 13$ 0.92
$25\pm 9$ $14.3\pm 2.2$ $5842\pm 25$ $72\pm 7$
6 $5\pm 5 11$$1.8\pm 1.3$ $5895\pm 16$ $50\pm 3$ $450\pm 280$ 0.84
$19\pm 10$ $12.8\pm 2.1$ $5853\pm 23$ $71\pm 7$
7 $3\pm 8 11$$1.0\pm 1.9$ $5908\pm 24$ $47\pm 5$ $420\pm 290$ 0.86
$52\pm 4$ $22.4\pm 1.2$ $5763\pm 8$ $72\pm 13$
8 $2.5\pm 1.6 11$$1.07\pm 0.29$ $5901.6\pm 2.1$ $49.6\pm 1.4$ $1000\pm 120$ 1.02
$9\pm 5$ $10.2\pm 0.6$ $5899\pm 5$ $56\pm 3$
9 $8\pm 3$ $9.8\pm 0.5$ $5931\pm 4$ $41.9\pm 2.2$ $140\pm 30$ 1.08
$1.3\pm 0.8$ $8.64\pm 0.06$ $5925.8\pm 1.0$ $48.6\pm 0.4$
10 $1.5\pm 0.08$ $8.41\pm 0.16$ $5946\pm 8$ $40.3\pm 2.6$ $2000\pm 600$ 0.97
0.0 (fixed) $8.3\pm 0.5$ $5933\pm 19$ $47\pm 6$
11 $1.95\pm 0.14$ $8.50\pm 0.16$ $5946.5\pm 1.8$ $39\pm 3$ $2100\pm 700$ 0.98
0.0 (fixed) $8.4\pm 0.4$ $5926.4\pm 2.7$ $50\pm 4$
12 $1.94\pm 0.13$ $8.50\pm 0.08$ $5946.5\pm 1.8$ $39.0\pm 1.9$ $2000\pm 700$ 1.01
0.0 (fixed) $8.4\pm 0.4$ $5926.5\pm 2.3$ $50\pm 4$
13 $2.6\pm 0.4$ $8.5\pm 0.4$ $5945.2\pm 2.3$ $43\pm 14$ $2500\pm 900$ 1.05
0.0 (fixed) $8.6\pm 0.9$ $5930\pm 10$ $42\pm 16$
14 0.0 (fixed) $8.5\pm 0.5$ 6.0 (fixed) $12\pm 6$ $1700\pm 600$ 0.82
–//– $8.5\pm 0.5$ –//– $15\pm 6$
15 0.0 (fixed) $9.7\pm 0.2$ 6.0 (fixed) 0.0 (fixed) 0.0 (fixed) 0.10
–//– $9.44\pm 0.05$ –//– –//–
As it is seen suggested approximation functions (4), (5) provide reasonable fit qualities at smallest $\sqrt{s_{\scriptsize{\mbox{min}},1}}=3$ GeV and, may be, at $\sqrt{s_{\scriptsize{\mbox{min}},2}}=5$ GeV; statistically acceptable $\chi^{2}/\mbox{n.d.f.}$ are obtained at any other $\sqrt{s_{\scriptsize{\mbox{min}}}}$. Thus, as previously Okorokov-IJMPA-25-5333-2010 ; Okorokov-IJMPA-32-1750175-2017 the parameterizations for the terms of the set $\mathcal{G}_{1}$ deduced within AQFT show good agreement with available experimental data. The recent RHIC PLB-808-135663-2020 and LHC EPJC-83-441-2023 ; EPJC-79-785-2019 results shown by open symbols in Fig. 1 agree well with general trends for global $pp$ scattering parameters (Fig. 1a, c) especially for $\sigma_{\scriptsize{\mbox{t}}}^{+}$ (Fig. 1a). Therefore, one can expect the new data with respect to the DB24 will negligibly effect on the values of fit parameters, especially for high $s_{\scriptsize{\mbox{min}}}$. This qualitative hypothesis is quite confirmed by many results of our previous studies Okorokov-CPC-46-083105-2022 ; Okorokov-IJMPA-25-5333-2010 ; Okorokov-IJMPA-32-1750175-2017 ; Okorokov-PAN-81-508-2018 made for earlier different versions of experimental database for terms of the set $\mathcal{G}_{1}$.
The values of all terms of the set $\mathcal{P}$ from Table 4 obtained for DB24 at low and intermediate $\bigl\{\sqrt{s_{\scriptsize{\mbox{min}},i}}\bigr\}_{i=1}^{7}$ mostly coincide with the values of corresponding fit parameters from Okorokov-IJMPA-25-5333-2010 ; Okorokov-IJMPA-32-1750175-2017 within errors for any $s_{\scriptsize{\mbox{min}}}$ indicated above. Fit qualities are (very) close to the published ones Okorokov-IJMPA-25-5333-2010 ; Okorokov-IJMPA-32-1750175-2017 at equal $\bigl\{\sqrt{s_{\scriptsize{\mbox{min}},i}}\bigr\}_{i=1}^{7}$. As consequence all features of the fit curves in Fig. 1 shown by solid ($\sqrt{s_{\scriptsize{\mbox{min}},2}}=5$ GeV) and dashed ($\sqrt{s_{\scriptsize{\mbox{min}},7}}=30$ GeV) lines agree with detailed discussion for our previous works Okorokov-IJMPA-25-5333-2010 ; Okorokov-IJMPA-32-1750175-2017 . Detailed analysis shows that values of some free parameters coincide with null or / and its asymptotic levels for modified Froissart–Martin limit within uncertainties starting from $\sqrt{s_{\scriptsize{\mbox{min}},10}}=60$ GeV already. In these cases corresponding experimental samples from DB24 are fitted at fixed values of such parameters (Table 2). Thus the functional behavior of $\sigma_{\scriptsize{\mbox{t}}}^{\pm}(s)$ approaches consecutively and gradually to the modified Froissart–Martin limit with growth of $\sqrt{s_{\scriptsize{\mbox{min}}}}$ starting with high value $\sqrt{s_{\scriptsize{\mbox{min}},14}}=1$ TeV. As it is seen from Table 4 the following relations are valid $\sigma_{\scriptsize{\mbox{t}}}^{\pm}(s)=\beta_{\pm}\sigma_{\scriptsize{\mbox{FM}}}$ and $\rho^{\pm}(s)=\rho_{\scriptsize{\mbox{FM}}}$ at highest for simultaneous fit $\sqrt{s_{\scriptsize{\mbox{min}},15}}=5$ TeV. Moreover $\sigma_{\scriptsize{\mbox{t}}}^{+}(s)=\sigma_{\scriptsize{\mbox{t}}}^{-}(s)$ due to coincidence of $\beta_{+}$ and $\beta_{-}$ within 1.04 standard deviation (Table 4). It means that any formulations of Pomeranchuk theorem are valid and all global scattering parameters considered here reach the modified Froissart–Martin limit at least in functional sense because of $\beta_{\pm}\ll\beta_{\scriptsize{\mbox{FM}}}$. The result $\sqrt{s_{a,\scriptsize{\mbox{FM}}}}=5$ TeV can be considered as qualitative indication only for the onset of asymptotic region for global scattering parameters with point of view of the theorems under consideration because of the reasons discussed above, namely, (i) limited data sample for $pp$ collisions, especially, for $\rho^{+}(s)$, (ii) availability of experimental results for $\rho^{+}(s)$ in (very) narrow energy range $\sqrt{s}=7-8$ TeV, and (iii) absence of $\bar{p}p$ data at all. All of these can influence on the robustness and stability for values of terms of the set $\mathcal{P}$ and, consequently, on physical conclusions. On the other hand, the result of the simultaneous fit at $\sqrt{s_{\scriptsize{\mbox{min}},15}}=5$ TeV agree qualitatively with the previous individual analysis of $\sigma_{\footnotesize\mbox{t}}^{+}(s)$ at $\sqrt{s_{\scriptsize{\mbox{min}},16}}=10$ TeV Okorokov-PAN-81-508-2018 provided the functional behavior of this quality close to the (modified) Froissart–Martin limit.
The large quantitative discrepancy between $\sigma_{\scriptsize{\mbox{t}}}^{\pm}$ and $\sigma_{\scriptsize{\mbox{FM}}}$ observed here and in previous studies Okorokov-IJMPA-25-5333-2010 ; Okorokov-IJMPA-32-1750175-2017 ; Okorokov-PAN-81-508-2018 even for multi-TeV region may indicate on the necessity of some novel dynamical mechnisms for $pp/\bar{p}p$ scattering which could be provide significant increase of total cross sections, i.e. $\beta_{\pm}$ values within AQFT approach. In general, Bose–Einstein condensation (BEC) can provide the noticeable growth of multiplicity of secondary pions in nucleus–nucleus collisions at sufficiently high energies Okorokov-PAN-87-172-2024 . Therefore, in general, BEC can affect the behavior of cross sections. Although this effect was not obtained for $pp$ up to $\sqrt{s}\lesssim 1$ PeV it can appears at larger energies. Then the BEC could be provide, for instance, the sharp increase of total cross section at certain $\sqrt{s}$ or in (very) narrow range of collision energy like to behavior of multiplicity of secondary pions Okorokov-PAN-87-172-2024 without violation of Froissart–Martin limit on future functional behavior of $\sigma_{\scriptsize{\mbox{t}}}(s)$. The scenario is similar to that for some parameter at 1st order phase transition with discontinuity and jump to the new (significantly higher) level of values. Thus, perhaps, BEC can be suggested as one of the possible dynamical mechanisms which may provide the approach to (modified) Froissart–Martin limit quantitatively in ”truly asymptotic region” which onset can be expected, in general, up to grand unified theory (GUT) energy domain in order of magnitude, i.e. $\sqrt{s_{a}}\gtrsim 10^{12}-10^{13}$ GeV Okorokov-IJMPA-33-1850077-2018 . This qualitative hypothesis request the future development and detailed study.
Fig. 2 shows the dependence of $\alpha_{\pm}$ (a), $\beta_{\pm}$ (b) and $\gamma_{\pm}$ (c) on the low boundary of fit range $\sqrt{s_{\scriptsize{\mbox{min}}}}$. Values of parameters $\alpha_{\pm}$ show sharp decrease at $\sqrt{s_{\scriptsize{\mbox{min}}}}\geq 30$ GeV approaching to the asymptotic level at $\sqrt{s_{\scriptsize{\mbox{min}}}}\geq 60$ GeV already (Fig. 2a) whereas the similar statement is valid for $\gamma_{\pm}$ at $\sqrt{s_{\scriptsize{\mbox{min}}}}\geq 1$ TeV (Fig. 2c). The parameter $\beta_{\pm}$ shows weaker dependence on $\sqrt{s_{\scriptsize{\mbox{min}}}}$ especially in the range $\sqrt{s_{\scriptsize{\mbox{min}}}}\gtrsim 50-60$ GeV (Fig. 2b). Some increase is observed for $\beta_{\pm}$ at these energies but $\beta_{\pm}/\beta_{\scriptsize{\mbox{FM}}}\lesssim 10^{-2}$ even at highest $\sqrt{s_{\scriptsize{\mbox{min}},15}}$ under consideration. The relations between free parameters of AQFT model agree with requests of Table 1 starting at $\sqrt{s_{\scriptsize{\mbox{min}},14}}=1$ TeV, i.e. AQFT allows the validity of both formulations of Pomeranchuk theorem at energies $\sqrt{s}\geq\sqrt{s_{a,\scriptsize{\mbox{P}}}}=1$ TeV.
The results obtained by simultaneous fitting at $\sqrt{s_{\scriptsize{\mbox{min}}}}\geq 1$ TeV are correct from the point of view of the methods of physical data analysis. However, it seems that aforementioned (very) limited sample of $\bar{p}p$ experimental data at $\sqrt{s}>1$ TeV and the complete absence of such data at $\sqrt{s}>2$ TeV allow only qualitative and preliminary conclusions about the presence of indications of the fulfillment of Pomeranchuk’s theorem at energies above 1 TeV and the achievement of the Froissart–Martin limiting behavior by the $\sigma_{\scriptsize{\mbox{t}}}^{\pm}$ in the functional sense at $\sqrt{s}\gtrsim 5$ TeV. Experimental data on $\bar{p}p$ scattering are very important in the multi-TeV region for more precise and definitive physical conclusions. However, at present, as it is known, projects to create a collider with $\bar{p}$ beam for energies comparable to the nominal energy of the LHC design, i.e. at $\sqrt{s}\sim 10$ TeV, are not being considered even in the medium-term perceptive in time. It seems that the development and implementation of a project for such an $\bar{p}p$ collider would be important from the point of view of the physics of strong interactions in general, and, in particular, for the study of global scattering parameters in $\bar{p}p$ interactions in a new energy region.
V Summary
The analytic parameterizations suggested within AQFT provide the quantitative description of energy dependence of global scattering parameters for $pp$ and $\bar{p}p$ scattering with robust values of fit parameters for recent DB24. The wide set of low boundary for fit domain is used. It allows the detailed scan of behavior of total cross sections and $\rho$-parameter in dependence of collision energy. Value of free parameters for AQFT model agree well with our previous works at low and intermediate $s_{\scriptsize{\mbox{min}}}$. Based on the present scan on $s_{\scriptsize{\mbox{min}}}$ extended to the high-energy region the numerical estimations are derived for the onsets of energy regions in which Pomeranchuk theorem and / or Froissart–Martin one is valid. Phenomenological parameterizations within AQFT and available experimental data allow the validity of both formulations of Pomeranchuk theorem at energies $\sqrt{s}\geq 1$ TeV. It is observed that all global scattering parameters considered here reach the modified Froissart–Martin limit at least in functional sense at $\sqrt{s}\geq 5$ TeV. But the numerical values of total cross sections are much smaller of corresponding asymptotic levels. Perhaps, the investigation of bosonic condensation will shed new light on the nature of this quantitative discrepancy. It seems that the development and building of new $\bar{p}p$ collider for study of multi-TeV energy region would be very useful, in particular, for physics of strong interaction processes at large distance and for direct search for the onset of asymptotic region for global scattering parameters.